Suppose f(x,y)= 4x2 + 3y2.

a. Calculate the partial derivatives of f.

b. Suppose f(x,y)= 16. Use the implicit function theorem to calculate dy/dx.

c. What is the value of dy/dx if x=1, y=2? 


1
Expert's answer
2021-11-16T10:37:58-0500

"f(x,y)=16"

Axis interception points of 16:

Y- intercept:(0,16)

Slope of 16: m=0.

Range: 16




(a)

"f(x,y)=4x^2+3y^2"

"\frac{\delta F}{\delta x}" is a partial derivative of F w.r.t. x.

"\frac{\delta}{\delta x}(4x^2+3y^2)=8x+0=8x"

Similarly,

"\frac{\delta}{\delta y}(4x^2+3y^2)=0+6y=6y"

(b)

At x=1 and y=2:

"F=4(1)^2+3(2)^2=16."

Total differential for F:

"=\frac{\delta F}{\delta x}.dx+\frac{\delta F}{\delta y}.dy"

"\delta F=8xdx+6ydy"

When "\delta F=0"

"\implies =8xdx-6ydy=0"

"8xdx=6ydy"

"\frac{dy}{dx}=\frac{8x}{6y}=\frac{4x}{3y}"

(c)

Value of "\frac{dy}{dx}" when x=1 and y=2:

"=\frac{4(1)}{3(2)}=\frac{4}{6}=\frac{2}{3}"